# TiM4Rec: An Efficient Sequential Recommendation Model Based on Time-Aware Structured State Space Duality Model

### v2

@@ 2.3. State Space Duality
From a computational efficiency standpoint, Gu et al. (SSD) innovatively transformed matrix $\bm{A}$ into a scalar value to facilitate SSM matrixization calculation and discovered that this variant of SSM exhibits duality with masked attention mechanisms. To preserve the linear complexity benefit of SSM in the sequence dimension, they leveraged the block properties of semi-separable matrices to implement linear attention computation, culminating in what they termed the Structured State Space Duality.
@@
$\bm{Y}$ $=SSM(\bm{A},\bm{B},\bm{C})(X)=\bm{M}\bm{X}$
$\bm{A}_{m:n}^{\times}$ $\coloneqq\prod_{k=m}^{n}\bm{A}_{k}=\bm{A}_{n}\bm{A}_{n-1}\cdots\bm{A}_{m}$
(3) $\bm{M}=\begin{bmatrix}\bm{C}_{0}^{\top}\bm{B}_{0}&&&&\\ \bm{C}_{1}^{\top}\bm{A}_{1}\bm{B}_{0}&\bm{C}_{1}^{\top}\bm{B}_{1}&&&\\ \bm{C}_{2}^{\top}\bm{A}_{2}\bm{A}_{1}\bm{B}_{0}&\bm{C}_{2}^{\top}\bm{A}_{2}\bm{B}_{1}&\bm{C}_{2}^{\top}\bm{B}_{2}&&\\ \vdots&\vdots&\ddots\\ \bm{C}_{i}^{\top}\bm{A}_{1:i}^{\times}\bm{B}_{0}&\bm{C}_{i}^{\top}\bm{A}_{2:i}^{\times}\bm{B}_{1}&\cdots&\bm{C}_{i}^{\top}\bm{B}_{j}\end{bmatrix}$
@@
(4) $\bm{L}=\begin{bmatrix}1\\ a_{1}&1\\ a_{2}a_{1}&a_{2}&1\\ \vdots&\vdots&\ddots&\ddots\\ a_{t-1}\ldots a_{1}&a_{t-1}\ldots a_{2}&\cdots&a_{t-1}&1\end{bmatrix}$
$\bm{M}_{i+1,j+1}=\bm{C}_{i}^{\top}\bm{A}_{i:j}\bm{B}_{j}\coloneqq\bm{C}_{i}^{\top}a_{i}\ldots a_{j+1}\bm{B}_{j}$
@@
$\bm{M}=\bm{L}\circ\bm{C}\bm{B}^{\top}$
$\bm{M}=\begin{bmatrix}\bm{C}_{0}^{\top}\bm{A}_{0:0}\bm{B}_{0}&&&&\\ \bm{C}_{1}^{\top}\bm{A}_{1:1}\bm{B}_{0}&\bm{C}_{1}^{\top}\bm{A}_{1:1}\bm{B}_{1}&&&\\ \bm{C}_{2}^{\top}\bm{A}_{2:1}\bm{B}_{0}&\bm{C}_{2}^{\top}\bm{A}_{2:1}\bm{B}_{1}&\bm{C}_{2}^{\top}\bm{A}_{2:2}\bm{B}_{2}&&\\ \vdots&\vdots&\ddots\\ \bm{C}_{i}^{\top}\bm{A}_{i:1}\bm{B}_{0}&\bm{C}_{i}^{\top}\bm{A}_{i:2}\bm{B}_{1}&\cdots&\bm{C}_{i}^{\top}\bm{A}_{i:i}\bm{B}_{i}\end{bmatrix}$
@@
$\bm{Y}=\bm{L}\circ\bm{C}\bm{B}^{\top}\bm{X}=\bm{L}\circ\bm{Q}\bm{K}^{\top}\bm{V}$
This concept is referred to as Duality in the SSD framework. By referring to the tensor contraction order in the calculation of linear attention (LinearAttention)—specifically using the associative property of matrix multiplication where $K^{T}V$ is computed first and subsequently left-multiplied by $Q$—the computational complexity is reduced from quadratic to linear in the sequence dimension without considering causal masks, a property that SSD also achieves.
$\bm{Y}=\left(\bm{Q}\bm{K}^{T}\right)\bm{V}\quad\Rightarrow\quad\bm{Y}=\bm{Q}\left(\bm{K}^{T}\bm{V}\right)$
$\bm{Y}=\left(\bm{C}\bm{B}^{T}\right)\bm{X}\quad\Rightarrow\quad\bm{Y}=\bm{C}\left(\bm{B}^{T}\bm{X}\right)$
To integrate causal masking without compromising linear computational complexity, the block properties of the 1-SS matrix $\bm{L}$ can be leveraged in conjunction with segment accumulation. This approach facilitates the implementation of causal linear attention, referred to as Structured Masked Attention (SMA) (SSD) in SSD. Since the SSD architecture is derived from the Mamba model, models based on the SSD architecture are referred to as Mamba2. Although SSD offers significant advantages in terms of computational load—particularly in tasks with high feature dimensions—the performance of SSD in sequential recommendation tasks is inferior compared to Mamba1. We provide a detailed analysis of this issue in the following sections.
@@ 3. METHOD
In this section, we offer a detailed introduction to our proposed framework, TiM4Rec. Following an overview of the overall structure of TiM4Rec, we delve into the reasons behind the performance degradation of SSD compared to SSM in low-dimensional scenarios within sequential recommendation tasks. Subsequently, we explain how our time-aware enhancement method, utilizing our proposed time-aware masked matrix with linear computational complexity, compensates for this loss.
3.1. Overview of TiM4Rec
As shown in Figure 2, the overall logic of TiM4Rec adheres to the description in section 2.1, However, an additional user interaction unix timestamp sequence $\mathcal{T}^{u_{k}}=\left\{t_{1},t_{2},\cdots,t_{L}\right\}\in\mathbb{R}^{L}$ is incorporated into the input data, with $\mathcal{T}^{u_{k}}$ corresponding one to one with the user interation sequence $\mathcal{S}^{u_{k}}$, where $L$ representing interaction sequence length. The sequence $\mathcal{S}^{u_{k}}$ can be transformed into an input sequence $\bm{X}\in\mathbb{R}^{T\times D}$ composed of dense vectors by the embedding table $\mathcal{E}$. Subsequently, a displacement subtraction is performed on sequence $\mathcal{T}^{u_{k}}$, and a zero is appended to the begining to generate the intearaction time difference sequence $\mathcal{D}^{u_{k}}=\left\{d_{1},d_{2},\cdots,d_{L}\right\}\in\mathbb{R}^{L}$, which be used in later stages. The Time-aware SSD Block enhances the model’s ability to capture potential user interests by integrating sequence $\mathcal{D}^{u_{k}}$ into a discretized SSD kernel. For the sake of convenience in the subsequent discussion, the superscript $u_{k}$ be omitted.
3.2. Time Difference Processing
As mentioned above, for the input interaction timestamp sequence $\mathcal{T}$, we transform it into an interaction time difference sequence $\mathcal{D}$ using the following method:
(5) $\mathcal{D}_{i}=\begin{cases}0,\quad&i=1\\ \mathcal{T}_{i}-\mathcal{T}_{i-1},\quad&i\in\left[2,L\right]\end{cases}$
@@ 3.2. Time Difference Processing
Subsequently, we apply dropout (srivastava2014dropout) and layer normalization (ba2016layer) to sequence $\mathcal{D}$ in order to eliminate data noise and optimize the data distribution for subsequent analysis.
3.3. Analyzing the Masked Matrix in SSD
@@
As discussed in section 2.3, the critical advancement from SSM to SSD resides in the scalarization of the transition matrix $\bm{A}$. When considering this modification from an attention perspective, the central aspect involves the alterations in the mask matric $\bm{L}$. Although the recursive operational constraints of matrix $\bm{A}$ in SSM prevent the direct and precise formulation of the masked matrix within eq. 3, it remains feasible to examine an equivalent expression for the final masked matrix:
$\bm{L}_{\text{SSM}}=\begin{bmatrix}I\\ \bm{A}_{1}&I\\ \bm{A}_{2}\bm{A}_{1}&\bm{A}_{2}&I\\ \vdots&\vdots&\ddots&\ddots\\ \bm{A}_{1:i}^{\times}&\bm{A}_{2:i}^{\times}&\cdots&\bm{A}_{i-1:i}^{\times}&I\end{bmatrix}$
@@
As illustrated in figure 3, in contrast to the masked matrix of SSD presented in eq. 4, the fundamental change in the masked matrix $L_{\text{SSM}}$ is that its mask coefficients are applied in a point-to-point manner on the individual item $\bm{E}_{i}=\left[e_{1},e_{2},\cdots,e_{n}\right]\in\mathbb{R}^{N}$ (The $\bm{A}$ matrix follows its definitions in SSM and SSD, respectively.):
@@
MSA stands for Vanilla Mask Self Attention Mechanism
However, owing to the semi-separable matrix structure inherent in the masked matrix of SSD, there is a pronounced attenuation effect on the interaction items that appear earlier in the sequence. This feature aligns SSD more closely with the actual user interaction patterns observed in real-world recommendation systems. Consequently, it is anticipated that, under equivalent conditions in low-dimensional settings, the SSM architecture exhibit superior performance in the realm of sequential recommendation, followed by SSD, with the Transformer architecture trailing behind. This hypothesis is further corroborated by the findings presented in our subsequent experimental section.
@@
Observing the differences between the equivalent mask matrices of SSD and SSM, it can be seen that the core of their mask matrices is recursively generated based on the state transition matrix $\bm{A}$. As described in section 2.3, the only difference is that SSD achieves more efficient matrix multiplication calculations by scalar quantization of matrix $\bm{A}$. Therefore, addressing the loss of the masked matrix caused by the scalar quantization of matrix $\bm{A}$ while maintaining the efficient matrix multiplication calculation of SSD has become the key to improvement.
3.4. Time-aware Masked Matrix
Let’s reconsider how the time difference matrix should be expressed in the context of attention mechanisms. The method proposed by TiSASRec (TiSASRec) is centered around acquiring the time position encoding for each item in the interaction sequence through interaction timestamps, ultimately leading to the generation of the corresponding attention bias matrix based on these encodings. Given that the Mamba architecture does not depend on positional encoding—as substantiated by the findings in Mamba4Rec (mamba4rec) and SSD4Rec (ssd4rec)—it is essential to directly explore the equivalent representation of the TiSASRec approach. Specifically, we should investigate how the resulting attention bias matrix encapsulates the temporal differences in product interactions.
For a sequence of interactive timestamps $\mathcal{T}$, we can derive the relative time difference matrix $\hat{\mathcal{D}}$ between each product through the following transformation:
(6) $\hat{\mathcal{D}}=\begin{bmatrix}\hat{d}_{11}\\ \mathbf{\underline{\hat{d}_{21}}}&\hat{d}_{22}\\ \hat{d}_{31}&\mathbf{\underline{\hat{d}_{32}}}&\hat{d}_{33}\\ \vdots&\vdots&\ddots&\ddots\\ \hat{d}_{L1}&\hat{d}_{L2}&\cdots&\mathbf{\underline{\hat{d}_{LL-1}}}&\hat{d}_{LL}\end{bmatrix},\quad\hat{d}_{mn}=t_{m}-t_{n}$
Directly applying matrix $\hat{\mathcal{D}}$ to the masked matrix $\bm{L}$ (as described in eq. 4) compromises the semi-separable matrix property of $\bm{L}$. This compromise prevents us from leveraging the efficient matrix computation techniques offered by the SSD architecture, consequently resulting in a return to quadratic computational complexity. Therefore, it is imperative to modify the insertion method of matrix $\hat{\mathcal{D}}$.

### v3

@@ 2.4. State Space Duality
From a computational efficiency standpoint, Dao et al. (SSD2024) innovatively transformed matrix $\bm{A}$ into a scalar value to facilitate SSM matrixization calculation and discovered that this variant of SSM exhibits duality with masked attention mechanisms. To preserve the linear complexity benefit of SSM in the sequence dimension, they leveraged the block properties of semi-separable matrices to implement linear attention computation, culminating in what they termed the Structured State Space Duality.
@@
(4) $\bm{Y}=SSM(\bm{A},\bm{B},\bm{C})(X)=\bm{M}\bm{X},$
(5) $\bm{A}_{m:n}^{\times}\coloneqq\prod_{k=m}^{n}\bm{A}_{k}=\bm{A}_{n}\bm{A}_{n-1}\cdots\bm{A}_{m},$
$\bm{M}$ $=\begin{bmatrix}\bm{C}_{0}^{\top}\bm{B}_{0}&&&&\\ \bm{C}_{1}^{\top}\bm{A}_{1}\bm{B}_{0}&\bm{C}_{1}^{\top}\bm{B}_{1}&&&\\ \bm{C}_{2}^{\top}\bm{A}_{2}\bm{A}_{1}\bm{B}_{0}&\bm{C}_{2}^{\top}\bm{A}_{2}\bm{B}_{1}&\bm{C}_{2}^{\top}\bm{B}_{2}&&\\ \vdots&\vdots&\ddots\\ \bm{C}_{i}^{\top}\bm{A}_{1:i}^{\times}\bm{B}_{0}&\bm{C}_{i}^{\top}\bm{A}_{2:i}^{\times}\bm{B}_{1}&\cdots&\bm{C}_{i}^{\top}\bm{B}_{j}\end{bmatrix},$
@@
(6) $\bm{L}=\begin{bmatrix}1\\ a_{1}&1\\ a_{2}a_{1}&a_{2}&1\\ \vdots&\vdots&\ddots&\ddots\\ a_{t-1}\ldots a_{1}&a_{t-1}\ldots a_{2}&\cdots&a_{t-1}&1\end{bmatrix},$
$\bm{M}_{i+1,j+1}=\bm{C}_{i}^{\top}\bm{A}_{i:j}\bm{B}_{j}\coloneqq\bm{C}_{i}^{\top}a_{i}\ldots a_{j+1}\bm{B}_{j},$
@@
(7) $\bm{M}=\bm{L}\circ\bm{C}\bm{B}^{\top},$
$\bm{M}=\begin{bmatrix}\bm{C}_{0}^{\top}\bm{A}_{0:0}\bm{B}_{0}&&&&\\ \bm{C}_{1}^{\top}\bm{A}_{1:1}\bm{B}_{0}&\bm{C}_{1}^{\top}\bm{A}_{1:1}\bm{B}_{1}&&&\\ \bm{C}_{2}^{\top}\bm{A}_{2:1}\bm{B}_{0}&\bm{C}_{2}^{\top}\bm{A}_{2:1}\bm{B}_{1}&\bm{C}_{2}^{\top}\bm{A}_{2:2}\bm{B}_{2}&&\\ \vdots&\vdots&\ddots\\ \bm{C}_{i}^{\top}\bm{A}_{i:1}\bm{B}_{0}&\bm{C}_{i}^{\top}\bm{A}_{i:2}\bm{B}_{1}&\cdots&\bm{C}_{i}^{\top}\bm{A}_{i:i}\bm{B}_{i}\end{bmatrix}.$
@@
(8) $\bm{Y}=\bm{L}\circ\bm{C}\bm{B}^{\top}\bm{X}=\bm{L}\circ\bm{Q}\bm{K}^{\top}\bm{V}.$
This concept is referred to as Duality in the SSD framework (SSD2024). By referring to the tensor contraction order in the calculation of linear attention (LinearAttention2020)—specifically using the associative property of matrix multiplication where $K^{T}V$ is computed first and subsequently left-multiplied by $Q$—the computational complexity is reduced from quadratic to linear in the sequence length without considering causal masks, a property that SSD also achieves.
$\bm{Y}=\left(\bm{Q}\bm{K}^{T}\right)\bm{V}\quad\Rightarrow\quad\bm{Y}=\bm{Q}\left(\bm{K}^{T}\bm{V}\right),$
$\bm{Y}=\left(\bm{C}\bm{B}^{T}\right)\bm{X}\quad\Rightarrow\quad\bm{Y}=\bm{C}\left(\bm{B}^{T}\bm{X}\right),$
To integrate causal masked matrix without compromising linear computational complexity, the block properties of the 1-SS matrix $\bm{L}$ can be leveraged in conjunction with segment accumulation. This approach facilitates the implementation of causal linear attention, referred to as Structured Masked Attention (SMA) (SSD2024) in SSD. Since the SSD architecture is derived from the Mamba model, models based on the SSD architecture are referred to as Mamba2. Although SSD offers significant advantages in terms of computational load—particularly in tasks with high feature dimensions—the performance of SSD in sequential recommendation tasks is inferior compared to Mamba1. We provide a detailed analysis of this issue in the following sections.
@@ 3. METHOD
In this section, we offer a detailed introduction to our proposed framework, TiM4Rec. Following an overview of the overall structure of TiM4Rec, we delve into the reasons behind the performance degradation of SSD compared to SSM in low-dimensional scenarios within sequential recommendation tasks. Subsequently, we explain how our time-aware enhancement method, utilizing our proposed Time-aware Masked Matrix with linear computational complexity, compensates for this performance degradation.
3.1. Problem Formulation and Method Overview
@@
Let $\mathcal{U}=\left\{u_{1},u_{2},\cdots,u_{|\mathcal{U}|}\right\}$ be the set of users and $\mathcal{V}=\left\{v_{1},v_{2},\cdots,v_{|\mathcal{V}|}\right\}$ be the set of items. For any user $u_{k}$, there exists an interaction sequence $\mathcal{S}^{u_{k}}=\left\{v_{1}^{u_{k}},v_{2}^{u_{k}},\cdots,v_{T}^{u_{k}}\right\}\in\mathbb{R}^{T}$ order by the user interaction unix timestamp sequence $\mathcal{T}^{u_{k}}=\left\{t_{1},t_{2},\cdots,t_{T}\right\}\in\mathbb{R}^{T}$, where $v_{i}^{u_{k}}$ represents the i-th item of user $u_{k}$ interaction and $T$ representing interaction sequence length. Sequential recommendation methodologies leverage the user’s historical interaction sequence $\mathcal{S}^{u_{k}}$ to forecast the next potential interaction item $v_{T+1}$ for the user $u_{k}$. Without loss of generality, we extract the preference representation $p^{u_{k}}$ for user $u_{k}$ from the historical interaction sequence $\mathcal{S}^{u_{k}}$ and the timestamp sequence $\mathcal{T}^{u_{k}}$, thereby predicting the probability distribution $\hat{\mathcal{Y}}\left(v_{T+1}|p^{u_{k}}\right)$ for the next potential interaction item. For convenience in the subsequent discussion, the superscript $u_{k}$ be omitted.
As illustrated in Fig. 2, the proposed TiM4Rec is a time-aware sequential recommendation model that leverages the SSD architecture through the Time-aware SSD blocks. The TiM4Rec model consists of multiple layers of Time-aware SSD Layers, each composed of a Time-aware SSD Block and a Feed Forward Network (FFN) (Transformer2017). We incorporate a special time difference processing method that allows the SSD kernel to integrate time difference information.
3.2. Data Preprocessing
3.2.1. Embedding Layer
In a manner analogous to previous models, we establish a learnable embedding table $\mathbb{E}=\left\{E_{1},E_{2},\cdots,E_{|\mathcal{V}|}\right\}\in\mathbb{R}^{|\mathcal{V}|\times D}$, the item IDs are converted from a sparse vector representation to a dense vector representation, where $E_{j}$ denotes the potential dense vector representation of the item $v_{j}$. The interaction sequence $\mathcal{S}$ can be transformed into an input sequence $\bm{X}\in\mathbb{R}^{T\times D}$ composed of dense vectors by the embedding table $\mathbb{E}$. Subsequently, we enhance the embedding table $\mathbb{E}$ representation learning process by employing dropout (Dropout2014) and layer normalization (LayerNorm2016) techniques. These methods mitigate overfitting and stabilize the training dynamics, thereby improving the robustness and generalization capabilities of the learned embeddings.
3.2.2. Time Difference Processing
To integrate the interactive timestamp sequence $\mathcal{T}$ into the SSD computation kernel and enhance the model’s ability to capture latent user interests, we perform differential processing on $\mathcal{T}$. The specific mechanisms of this integration will be discussed shortly. The interactive timestamp sequence $\mathcal{T}$ undergoes a displacement subtraction, with an initial zero appended, resulting in the time difference sequence $\mathcal{D}=\left\{d_{0},d_{1},\cdots,d_{T-1}\right\}\in\mathbb{R}^{T}$. The computation process is as follows:
(9) $\mathcal{D}_{i}=\begin{cases}0,\quad&i=0\\ \mathcal{T}_{i+1}-\mathcal{T}_{i},\quad&i\in\left[1,T-1\right]\end{cases}.$
@@ 3.2.2. Time Difference Processing
Subsequently, we apply dropout (Dropout2014) and layer normalization (LayerNorm2016) to sequence $\mathcal{D}$ in order to eliminate data noise and optimize the data distribution for subsequent analysis.
3.3. Time-Aware Linear Complexity Enhancement
3.3.1. Analyzing the Masked Matrix in SSD
@@
As discussed in section 2.4, the critical advancement from SSM to SSD resides in the scalarization of the transition matrix $\bm{A}$. When considering this modification from an attention perspective, the central aspect involves the alterations in the masked matrix $\bm{L}$. Although the recursive operational constraints of matrix $\bm{A}$ in SSM prevent the direct and precise formulation of the masked matrix within Eq. 5, it remains feasible to examine an equivalent expression for the final masked matrix:
(10) $\bm{L}_{\text{SSM}}=\begin{bmatrix}I\\ \bm{A}_{1}&I\\ \bm{A}_{2}\bm{A}_{1}&\bm{A}_{2}&I\\ \vdots&\vdots&\ddots&\ddots\\ \bm{A}_{1:i}^{\times}&\bm{A}_{2:i}^{\times}&\cdots&\bm{A}_{i-1:i}^{\times}&I\end{bmatrix},$
@@
As illustrated in Fig. 3, in contrast to the masked matrix of SSD presented in Eq. 6, the fundamental change in the masked matrix $L_{\text{SSM}}$ is that its mask coefficients are applied in a point-to-point manner on the individual item $\bm{E}_{i}=\left[e_{1},e_{2},\cdots,e_{n}\right]\in\mathbb{R}^{D}$ (The $\bm{A}$ matrix follows its definitions in SSM and SSD, respectively.):
@@
MSA stands for Vanilla Masked Self Attention Mechanism
However, owing to the 1-SS matrix structure inherent in the masked matrix of SSD, there is a pronounced attenuation effect on the interaction items that appear earlier in the sequence. This feature aligns SSD more closely with the actual user interaction patterns observed in real-world recommendation systems. Consequently, it is anticipated that, under equivalent conditions in low-dimensional settings, the SSM architecture exhibit superior performance in the realm of sequential recommendation, followed by SSD, with the Transformer architecture trailing behind. This hypothesis is further corroborated by the findings presented in our subsequent experimental section.
@@
Observing the differences between the equivalent masked matrices of SSD and SSM, it can be seen that the core of their mask matrices is recursively generated based on the state transition matrix $\bm{A}$. As described in section 2.4, the only difference is that SSD achieves more efficient matrix multiplication calculations by scalar quantization of matrix $\bm{A}$. Therefore, addressing the loss of the masked matrix caused by the scalar quantization of matrix $\bm{A}$ while maintaining the efficient matrix multiplication calculation of SSD has become the key to improvement.
3.3.2. Time-aware Structured Masked Matrix
Let’s reconsider the time-aware enhancement approach of TiSASRec (TiSASRec2020). The essence of this method lies in embedding the time difference matrix and processing it through the attention mechanism to generate a time-aware attention score matrix, which is then integrated into the final attention score matrix. If we regard the generated time-aware attention score matrix as a special type of masked matrix, it aligns with our previous analysis: employing a more refined mask matrix to extract potential user interest features can lead to improved recommendation performance. However, in the TiSASRec method, the computational steps involved in generating the time-aware masked matrix by embedding the time difference matrix and processing it through the attention mechanism incur substantial computational overhead. Therefore, it is imperative to identify a direct and efficient approach to convert the time difference matrix into a time-aware masked matrix.
For a sequence of interactive timestamps $\mathcal{T}$ (The definition is the same as in section 3.1), we can derive the relative time difference matrix $\hat{\mathcal{D}}$ between each product through the following transformation:
(11) $\hat{\mathcal{D}}=\begin{bmatrix}\hat{d}_{11}\\ \mathbf{\underline{\hat{d}_{21}}}&\hat{d}_{22}\\ \hat{d}_{31}&\mathbf{\underline{\hat{d}_{32}}}&\hat{d}_{33}\\ \vdots&\vdots&\ddots&\ddots\\ \hat{d}_{T1}&\hat{d}_{T2}&\cdots&\mathbf{\underline{\hat{d}_{TT-1}}}&\hat{d}_{TT}\end{bmatrix},\quad\hat{d}_{mn}=t_{m}-t_{n},$
Directly applying matrix $\hat{\mathcal{D}}$ to the masked matrix $\bm{L}$ (as described in Eq. 6) compromises the 1-SS matrix property of $\bm{L}$. This compromise prevents us from leveraging the efficient matrix computation techniques offered by the SSD architecture, consequently resulting in a return to quadratic computational complexity. Therefore, it is imperative to modify the insertion method of matrix $\hat{\mathcal{D}}$.